# General-relativistic magnetar magnetospheres in 3D with physics-informed neural networks

### v1

@@ Abstract
Magnetar phenomena are likely intertwined with the location and structure of magnetospheric currents. While general-relativistic effects may be important in shaping the force-free equilibria describing static configurations, most studies have been restricted to axial symmetry. Using a novel methodology based on physics-informed neural networks, fully three-dimensional configurations of varying stellar compactness are constructed. Realistic profiles for surface currents, qualitatively capturing the geometry of observed hotspots, are applied as boundary conditions to deduce the amount of free energy available to fuel outburst activity. It is found that the lowest-energy solution branches permit only a $\approx 30\%$ excess relative to current-starved solutions in axisymmetric cases with global twists, regardless of compactness, reducing to $\approx 5\%$ in 3D models with localised spots. Accounting for redshift reductions to their inferred dipole moments from timing data, explaining magnetar burst energetics therefore becomes more difficult unless the field hosts non-negligible multipoles. Discussions on other aspects of magnetar phenomena are also provided.
@@
pubyear: 2025pagerange: General-relativistic magnetar magnetospheres in 3D with physics-informed neural networksB.1
@@ 1 Introduction
Magnetars are a class of young, highly magnetised, and slowly rotating neutron stars (Thompson & Duncan, 1995, 1996). They are observed mainly in X-rays, either through unpredictable, explosive transients (short bursts, outbursts, giant flares) or through persistent, pulsed thermal emissions (Turolla et al., 2015; Kaspi & Beloborodov, 2017). Both types of activity are thought to be linked to twisted, magnetospheric substructure (Thompson et al., 2002; Beloborodov & Thompson, 2007). These currents are thought to be sourced by the displacement of field-line footpoints instigated by mechanical stresses applied to the crust. Stress buildup can take place either abruptly, via starquakes (e.g. Franco et al., 2000), or gradually, due to elastic/plastic motions (e.g. Lander & Gourgouliatos, 2019), such that a magnetospheric instability arises once a critical threshold is breached. Observational (Tiengo et al., 2013; Younes et al., 2022) and theoretical (Beloborodov, 2009; Perna & Pons, 2011) evidence suggests that twists develop in localised regions, forming magnetospheric loops akin those in the solar corona.
@@
Axisymmetric, force-free (FF) ‘Grad-Shafranov’ equilibria have been studied extensively in the neutron-star context (e.g. Pili et al., 2015; Akgün et al., 2017). Recently, a more general treatment employing a full three-dimensional setup was presented in Stefanou et al. (2023a) (see also Mahlmann et al., 2023). These works establish the basic features in the modelling of magnetar magnetospheres with an important consensus being that, depending on the boundary conditions imposed at the stellar surface and on the toroidal field in particular, it is possible that there is a unique solution, multiple ones, or none at all (see Aly, 1984). In cases where multiple solutions exist, higher-energy equilibria host disconnected magnetic domains (‘islands’) and it is expected that only the lowest-energy branch is stable (Yu, 2012; Mahlmann et al., 2019). In other words, some form of rearrangement must take place if a magnetospheric configuration becomes ‘overtwisted’ such that a stable state is restored after plasmoid ejection. The energy released in these events is channelled to particles and then radiated away, potentially feeding a magnetar outburst (Lyutikov, 2003, 2006). In order to explain the observational phenomenology, it is thus crucial to understand what factors could adjust the available energy budget in a twisted magnetosphere.
@@
For compact objects, general-relativistic (GR) effects reshape the solution landscape and a self-consistent treatment should take them into account. Notably, Kojima (2017) reports that the energy that can be stored – relative to the untwisted, potential state – in a GR magnetosphere can be several times larger than the corresponding flat-space maximum for ultracompact stars (see also Yu, 2011; Kojima & Okamoto, 2018). However, their most energetic models also correspond to (likely) unstable branches containing disconnected magnetic domains, as described above. Nevertheless, even in the more lenient case of lowest-energy configurations, the stored energy surpasses that of flat-space equilibria for a fixed dipole moment because the surface field is augmented by GR factors (e.g. Petterson, 1974). On the other hand, the spin-down luminosity scales with stellar compactness for the same reason (Rezzolla & Ahmedov, 2004). From an observational perspective, this means that the surface field, as inferred from timing data, is lower than the Newtonian case and thus even if more energy can be stored in a relative sense, the absolute value available for outburst activity may be lower in reality.
@@ 2.1 Problem setup
Deformations due to rotation and hydromagnetic forces can be safely neglected in the metric surrounding a slow star with a sub-virial magnetic field ($B\ll 10^{18}$ G). The spacetime line element for cases of interest is therefore well-described by the Schwarzschild solution (Rädler et al., 2001; Rezzolla & Ahmedov, 2004; Kojima, 2017)
@@
In the magnetosphere of a magnetar, the magnetic force is dominant and all other forces can be safely neglected because of the rarity of magnetospheric plasma, the low mass of the charge carriers ($e^{\pm}$), and the slow rotation. This is the well-known FF regime, which is mathematically described by
×$\bm{\nabla}\times(e^{\nu}\bm{B})=\alpha\bm{B},$ (2)
@@
where $\mu_{B}$ is the magnetic dipole moment (as measured by a distant observer) and the subscript * denotes a restriction to the stellar surface. The scaling of expression (5), which entails an effective boost to the magnitude of the magnetic field in the GR case, is depicted in Figure 1. At fixed $\mu_{B}$, an increase of $\gtrsim 65\%$ can be seen for very compact stars with $M/R\gtrsim 0.25$, which impacts not only the magnetospheric structure (Sec. 4) but also the spindown luminosity (Sec. 5). In the absence of magnetospheric currents (i.e. when $\alpha=0$), equation (2) with boundary condition (5) can be solved analytically to give the potential solution (e.g. Petterson, 1974)
@@ 3 Method: PINNs
For a set of coordinates $\bm{x}$ is some domain $\mathcal{D}$, a general PDE can be written as
@@
We impose boundary conditions by hard-enforcement (Dong & Ni, 2021; Sukumar & Srivastava, 2022; Xiao et al., 2024; Urbán et al., 2025). This means that we do not define $B_{r},B_{\theta},B_{\phi},\alpha$ to be the outputs of the neural network directly. Instead, we modify the outputs so that they automatically satisfy the desired boundary conditions before inserting them in the equations to construct $\mathcal{L}$ (see Appendix A).
@@ 4 Results
Figure 2: Isocontours of the twist density $\alpha$ in the $x-z$ plane for axisymmetric solutions for a compact star with $M/R=0.25$ and current particulars $s=0.243$ and $m=3$ (brighter shades indicate greater values of $\alpha$). (a): The low energy solution. (b) The high energy solution, where a detached region (‘island’) protrudes from the equator.
@@ 4.1 Axisymmetric models
We focus first on axisymmetric models. We produce a sequence of magnetospheric solutions for fixed $m=3$ and various values of the parameter $s$ in equation (8). Figure 2(a) shows an example case with $s=0.243$ for a compact star with $M/R=0.25$. Field lines of fixed $\alpha$ are plotted in the $x-z$ plane. We see that the magnetospheric configuration is that of an ‘inflated dipole’, as is characteristic of such ‘power-law’ models from the literature (e.g., Flyer et al., 2004; Pili et al., 2015; Akgün et al., 2018). In a more direct sense, the field structure can be compared with the middle panel in figure 3 from Kojima (2017) which also has $M/R=0.25$ and exhibits the same qualitative features: field lines are pinched near the equator and inflated at larger radii by toroidal pressures, encapsulated by the twist density (reaching a maximum of $\approx 4R^{-1}$ near the surface).
This pinching effect becomes more pronounced as the parameter $s$ increases, leading to a regime where the elliptic equation (2) in its axisymmetric limit admits degenerate solutions with a fully detached ‘island’. These solutions have the same boundary conditions (i.e. $s$ value) and compactness but differ in energy and magnetospheric structure. One rather extreme case is shown in Figure 2(b), where a magnetically-disconnected domain is visible around $r\sim 2R$. We find that the relative energy excess in this case is $E^{(0.25)}_{e}=1.28$, which can be compared with the ‘stable’ case shown in the left panel which has a free energy that is a factor 60 smaller, $E^{(0.25)}_{e}=0.02$. We use the word stable since Mahlmann et al. (2019) showed that such ‘upper branch’ solutions with islands are dynamically unstable and, therefore, unlikely to be reached in astrophysically relevant scenarios. While islands are not special to GR cases, this huge energy disparity highlights the magnetic confinement effect discussed by Kojima (2017). For any compactness, increasing (decreasing) the value of $s$ tends to shift the island further (closer) away from the star, until eventually no solution exists at all.
@@
Figure 3 shows the dependence of the excess energy, $E_{e}$, on the maximum value of the twist density at the surface, $\alpha_{*}^{\mathrm{max}}$ (which depends both on $s$ and $M/R$), for different values of the compactness. For values of $\alpha_{*}^{\mathrm{max}}\gtrsim 4R^{-1}$, we find that no solutions exist for any compactness. The sequence has a turning point at this value, bifurcating into two branches with different energies. Physically, the critical point $\alpha_{*}^{\mathrm{max}}\approx 4R^{-1}$ corresponds to the most energetic, stable configuration. Notably, the excess is largely independent of the compactness and the energy stored in the magnetosphere is $\approx 30\%$ higher than the untwisted case at the turning point. Such a value agrees with Akgün et al. (2018), who found $E_{e}^{\rm max}\approx 25\%$ in the Newtonian limit. This indicates that, for a fixed dipole moment $\mu_{B}$, GR magnetospheres can store more energy than their Newtonian counterparts in absolute terms because $E^{(M)}_{0}$ increases monotonically with $M/R$ (though see Sec. 5). In a relative sense however, the free energy that stable branches can hold is practically independently of compactness.
@@
To understand the impact of the stellar compactness, we study three families of models with $M/R=0,0.1,0.25$. Example configurations for two different values of $\alpha_{*}^{\rm max}$ (which coincides with $\alpha_{0}$ in this case, see equation 9) are shown in Figure 4, corresponding to either a mild ($\alpha_{*}^{\rm max}=1.0R^{-1}$; top) or extreme ($\alpha_{*}^{\rm max}=2.5R^{-1}$; bottom panels) twist with increasing compactness going from left to right. We see that, much like in the axisymmetric case, the toroidal component of the field exerts a magnetic pressure which ‘inflates’ the magnetosphere in the current-filled region for larger $\alpha_{*}^{\rm max}$ noting that the scale extends to $r\sim 4\ R$. The size of the closed, bulbous region protruding from the spots decreases monotonically as a function of compactness (by a factor $\sim 1.5$ comparing the $M/R=0.25$ and $M=0$ cases), which further softens the effect of the twist. Because the field is locally stronger near the surface in cases with higher compactness (equation 5), the magnetic pressure is larger for a given $\mu_{B}$ which may be responsible for this ‘shrinking’ effect. In all cases, we see that only the field lines that emerge from the hotspots are twisted, which can be contrasted with the axisymmetric case where twists are always ‘global’ in some sense. Larger values of $\alpha_{*}^{\rm max}$ lead to more twisted configurations as expected, since the currents are stronger in this case. For example, comparing the final column we see a ‘weaving’ pattern only in in the larger twist case and that the toroidal blowout is a factor $\lesssim 2$ more effective.
@@
In Figure 5 we show the relative excess energy $E_{e}$ as a function of $\alpha_{*}^{\mathrm{max}}$. Interestingly, neither the functional dependence of the energy on $\alpha_{*}^{\rm max}$ nor the maximum achievable twist seem to be affected by the compactness. In particular, though a one-to-one comparison is made difficult by the different choices of current function, these features were also observed in stable axisymmetric cases (Fig. 3). Quantitatively however, because the twist is more localised the excess energy is lower. Note, however, that models with higher $M/R$ have higher absolute energy for a fixed dipole moment because the surface field is effectively boosted (Fig. 1).
@@ 5.1 Spindown luminosity and surface-field inferences
where $\Omega_{*}$ is the star-frame velocity. These facts together imply that the spindown luminosity increases significantly as a function of compactness, as established by Rezzolla & Ahmedov (2004).
If one includes corrections due to magnetospheric drag (Pétri, 2016), matching the rotational kinetic-energy loss to the (untwisted) spindown luminosity produces an expression for the polar field strength (restoring units)
$B_{p}^{2}=\left(1-\frac{2GM}{c^{2}R}\right)^{4}\frac{3c^{3}P\dot{P}I_{0}}{2\pi^{2}R^{6}\left(a+b\sin^{2}\chi\right)},$ (15)
which one may recognize as the Spitkovsky (2006) formula multiplied by a redshift prefactor for (GR-corrected) moment of inertia $I_{0}$ and magnetic inclination angle $\chi$. The factors $a$ and $b$ are fitting coefficients of order unity; in fact they are also generally larger than their Newtonian counterparts (see Table 2 in Pétri, 2016) indicating a further reduction. For slow magnetars we take $a=1$ and $b=1.5$.
@@
$B^{\rm GR}_{p}\sim 1.4\times 10^{19}\sqrt{P\dot{P}}\text{ G},$ (16)
@@
Figure 6 displays some relevant dimensionless ratios assuming the factors $a$ and $b$ in expression (15) are independent of compactness for simplicity. The top (dotted) and bottom (dashed) lines display the GR-to-Newtonian magnetospheric energy ratios for either a fixed polar field strength or one that decreases as a function of compactness (solid), respectively. The latter is more appropriate for a timing-inferred polar strength. We see that while the energy itself scales with the compactness (see Appendix B), it does so at a rate that is slower than the Poynting flux and hence one may anticipate magnetar magnetospheres to house less energy in the GR case if the polar field is not held fixed but rather deduced from $P-\dot{P}$ data. Since the available free energy is roughly constant as a function of compactness (see Fig. 3), a decrease by a factor $\gtrsim 3$ in the inferred (via timing) $E_{e}$ could be expected for a typical neutron star ($M/R\gtrsim 0.17$).
Figure 6: Ratios of dipolar magnetospheric energies normalised by the Newtonian values using the (GR or Newtonian) Tolman-VII equation of state as a function of compactness, with either fixed (dotted) or decreasing (dashed) polar field strengths. The latter applies to cases where the field is inferred from timing data, the GR-to-Newtonian ratio of which is depicted by the solid line.

### v2

@@ Abstract
Magnetar phenomena are likely intertwined with the location and structure of magnetospheric currents. General-relativistic effects are important in shaping the force-free equilibria describing static configurations, though most studies have quantified their impact only in cases of axial symmetry. Using a novel methodology based on physics-informed neural networks, fully three-dimensional configurations of varying stellar compactness are constructed. Realistic profiles for surface currents, qualitatively capturing the geometry of observed hotspots, are applied as boundary conditions to deduce the amount of free energy available to fuel outburst activity. It is found that the lowest-energy solution branches permit only a $\approx 30\%$ excess relative to current-starved solutions in axisymmetric cases with global twists, regardless of compactness, reducing to $\approx 5\%$ in 3D models with localised spots. Accounting for redshift reductions to their inferred dipole moments from timing data, explaining magnetar burst energetics therefore becomes more difficult unless the field hosts non-negligible multipoles. Discussions on other aspects of magnetar phenomena are also provided.
@@
pubyear: 2025pagerange: General-relativistic magnetar magnetospheres in 3D with physics-informed neural networks12
@@ 1 Introduction
Magnetars are a class of young, highly magnetised, and slowly rotating neutron stars (Thompson & Duncan, 1995, 1996). They are observed mainly in X-rays, either through unpredictable, explosive transients (short bursts, outbursts, giant flares) or through persistent, pulsed thermal emissions (Turolla et al., 2015; Kaspi & Beloborodov, 2017). Both types of activity are thought to be linked to twisted, magnetospheric substructure (Thompson et al., 2002; Beloborodov & Thompson, 2007). These currents are thought to be sourced by the displacement of field-line footpoints instigated by mechanical stresses applied to the crust. Stress accumulation is typically considered a gradual process driven by the interior field evolution, continuing until a critical threshold is breached. At this point, stresses are rapidly released (typically on much faster timescales) via abrupt starquakes (e.g. Perna & Pons, 2011; Pons & Perna, 2011; Franco et al., 2000), plastic motions (e.g. Lander & Gourgouliatos, 2019), or magnetospheric instabilities (Carrasco et al., 2019). Observational (Tiengo et al., 2013; Younes et al., 2022) and theoretical (Beloborodov, 2009) evidence suggests that twists develop in localised regions, forming magnetospheric loops akin those in the solar corona.
@@
Axisymmetric, force-free (FF) ‘Grad-Shafranov’ equilibria have been studied extensively in the neutron-star context (e.g. Pili et al., 2015; Akgün et al., 2017). Recently, a more general treatment employing a full three-dimensional setup was presented in Stefanou et al. (2023a) (see also Carrasco et al., 2019; Mahlmann et al., 2023). These works establish the basic features in the modelling of magnetar magnetospheres with an important consensus being that, depending on the boundary conditions imposed at the stellar surface and on the toroidal field in particular, it is possible that there is a unique solution, multiple ones, or none at all (see Aly, 1984). In cases where multiple solutions exist, higher-energy equilibria host disconnected magnetic domains (‘islands’) and it is expected that only the lowest-energy branch is stable (Yu, 2012; Mahlmann et al., 2019). In other words, some form of rearrangement must take place if a magnetospheric configuration becomes ‘overtwisted’ such that a stable state is restored after plasmoid ejection. The energy released in these events is channelled to particles and then radiated away, potentially feeding a magnetar outburst (Lyutikov, 2003, 2006; Parfrey et al., 2013; Sharma et al., 2023). In order to explain the observational phenomenology, it is thus crucial to understand what factors could adjust the available energy budget in a twisted magnetosphere.
@@
For compact objects, general-relativistic (GR) effects reshape the solution landscape and a self-consistent treatment should take them into account. Focussing on the GR effects on the balance of magnetic stresses, Kojima (2017) reports that the energy that can be stored – relative to the untwisted, potential state – in a GR magnetosphere can be several times larger than the corresponding flat-space maximum for ultracompact stars (see also Yu, 2011; Kojima & Okamoto, 2018). However, their most energetic models also correspond to (likely) unstable branches containing disconnected magnetic domains, as described above. Nevertheless, even in the more lenient case of lowest-energy configurations, the stored energy surpasses that of flat-space equilibria for a fixed dipole moment because the surface field is augmented by GR factors (e.g. Petterson, 1974). On the other hand, the spin-down luminosity scales with stellar compactness for the same reason (Rezzolla & Ahmedov, 2004). From an observational perspective, this means that the surface field, as inferred from timing data, is lower than the Newtonian case and thus even if more energy can be stored in a relative sense, the absolute value available for outburst activity may be lower in reality.
@@ 2.1 Problem setup
Deformations due to rotation and hydromagnetic forces can be safely neglected in the metric surrounding a slow [^1] Even for a period of $P\lesssim 2$ ms, the $tt$-component of the metric changes by at most $\sim 10\%$ for compact stars with realistic equations of state, as verified by the RNS code (Stergioulas & Friedman, 1995). star with a sub-virial magnetic field ($B\ll 10^{18}$ G). The spacetime line element for cases of interest is therefore well-described by the Schwarzschild solution (Rädler et al., 2001; Rezzolla & Ahmedov, 2004; Kojima, 2017)
@@
In the magnetosphere of a magnetar, the magnetic force overwhelmingly dominates, allowing all other forces to be neglected: the magnetic pressure far exceeds the plasma pressure, the $e^{\pm}$ charge carriers carry negligible momentum, inertia is irrelevant, and rotation is too slow to play a significant role. This is the well-known FF regime, which is mathematically described by
$\bm{\nabla}\times(e^{\nu}\bm{B})=\alpha\bm{B},$ (2)
@@
where $\mu_{B}$ is the magnetic dipole moment (as measured by a distant observer) and the subscript * denotes a restriction to the stellar surface. The scaling of expression (5), which entails an effective amplification to the magnitude of the magnetic field in the GR case, is depicted in Figure 1. At fixed $\mu_{B}$, an increase of $\gtrsim 65\%$ can be seen for very compact stars with $M/R\gtrsim 0.25$, which impacts not only the magnetospheric structure (Sec. 4) but also the spindown luminosity (Sec. 5). In the absence of magnetospheric currents (i.e. when $\alpha=0$), equation (2) with boundary condition (5) can be solved analytically to give the potential solution (e.g. Petterson, 1974)
@@ 3 Method: PINNs
For a set of coordinates $\bm{x}$ in some domain $\mathcal{D}$, a general PDE can be written as
@@
We impose boundary conditions by hard-enforcement (Dong & Ni, 2021; Sukumar & Srivastava, 2022; Xiao et al., 2024; Urbán et al., 2025). This means that the desired boundary conditions are implemented on the outputs of the network instead of the loss function and are exactly satisfied by construction (see Appendix A).
@@ 4 Results
Figure 2: Isocontours of the twist density $\alpha$ in the xzx-z plane for axisymmetric solutions for a compact star with $M/R=0.25$, $s=0.243$ and $m=3$ (brighter shades indicate greater values of $\alpha$). (a): The low energy solution. (b) The high energy solution, where a detached region (‘island’) protrudes from the equator. Panel (c) shows a 3D view of the high energy solution, where a disconnected field line with $\alpha\simeq 3.5R^{-1}$ circles the star. Other twisted lines connected to the surface of the star are also shown for comparison.
@@ 4.1 Axisymmetric models
We focus first on axisymmetric models. We produce a sequence of magnetospheric solutions for fixed $m=3$ and various values of the parameter $s$ in equation (8). Figure 2(a) shows an example case with $s=0.243$ for a compact star with $M/R=0.25$. Field lines of fixed $\alpha$ are plotted in the xzx-z plane. We see that the magnetospheric configuration is that of an ‘inflated dipole’, as is characteristic of such ‘power-law’ models from the literature (e.g., Flyer et al., 2004; Pili et al., 2015; Akgün et al., 2018). In a more direct sense, the field structure can be compared with the middle panel in figure 3 from Kojima (2017) which also has $M/R=0.25$ and exhibits the same qualitative features: field lines are pinched near the equator and inflated at larger radii by toroidal pressures, encapsulated by the twist density (reaching a maximum of $\approx 4R^{-1}$ near the surface).
This pinching effect becomes more pronounced as the parameter $s$ increases, leading to a regime where the elliptic equation (2) in its axisymmetric limit admits degenerate solutions with a fully detached ‘island’. These solutions have the same boundary conditions (i.e. $s$ value) and compactness but differ in energy and magnetospheric structure. One rather extreme case is shown in Figure 2(c), where a magnetically-disconnected domain is visible around $r\sim 2R$. We find that the relative energy excess in this case is $E^{(0.25)}_{e}=1.28$, which can be compared with the ‘stable’ case shown in the left panel which has a free energy that is a factor 60 smaller, $E^{(0.25)}_{e}=0.02$. We use the word stable since Mahlmann et al. (2019) showed that such ‘upper branch’ solutions with islands are dynamically unstable and, therefore, unlikely to be reached in astrophysically relevant scenarios. While islands are not special to GR cases, this huge energy disparity highlights the magnetic confinement effect discussed by Kojima (2017). For any compactness, increasing the value of $s$ tends to shift the island further away from the star until eventually no solution exists at all, and vice versa.
@@
Figure 3 shows the dependence of the excess energy, $E_{e}$, on the maximum value of the twist density at the surface, $\alpha_{*}^{\mathrm{max}}$ (which depends both on $s$ and $M/R$), for different values of the compactness. For values of $\alpha_{*}^{\mathrm{max}}\gtrsim 4R^{-1}$, we find that no solutions exist for any compactness. The sequence has a turning point at this value, bifurcating into two branches with different energies. In order to facilitate convergence and build sequences of degenerate models, we initialise the solver using the previous converged solution with slightly modified values of the parameters defining the boundary condition. Physically, the critical point $\alpha_{*}^{\mathrm{max}}\approx 4R^{-1}$ corresponds to the most energetic, stable configuration. Notably, the excess is largely independent of the compactness and the energy stored in the magnetosphere is $\approx 30\%$ higher than the untwisted case at the turning point. Such a value agrees with Akgün et al. (2018), who found $E_{e}^{\rm max}\approx 25\%$ in the Newtonian limit. This indicates that, for a fixed dipole moment $\mu_{B}$, GR magnetospheres can store more energy than their Newtonian counterparts in absolute terms because $E^{(M)}_{0}$ increases monotonically with $M/R$ (though see Sec. 5). In a relative sense however, the free energy that stable branches can hold is practically independently of compactness.
@@
To understand the impact of the stellar compactness, we study three families of models with $M/R=0,0.1,0.25$. Example configurations for two different values of $\alpha_{*}^{\rm max}$ (which coincides with $\alpha_{0}$ in this case, see equation 9) are shown in Figure 4, corresponding to either a mild ($\alpha_{*}^{\rm max}=1.0R^{-1}$; top) or extreme ($\alpha_{*}^{\rm max}=2.5R^{-1}$; bottom panels) twist with increasing compactness going from left to right. We see that, much like in the axisymmetric case, the toroidal component of the field exerts a magnetic pressure which ‘inflates’ the magnetosphere in the current-filled region for larger $\alpha_{*}^{\rm max}$ noting that the scale extends to $r\sim 4\ R$. The equatorial extent of the twisted flux bundle protruding from the spots decreases monotonically as a function of compactness (by a factor $\sim 1.5$ comparing the $M/R=0.25$ and $M=0$ cases), which further softens the effect of the twist. Because the field is locally stronger near the surface in cases with higher compactness (equation 5), the magnetic pressure is larger for a given $\mu_{B}$ which may be responsible for this ‘shrinking’ effect. In all cases, we see that only the field lines that emerge from the hotspots are twisted, which can be contrasted with the axisymmetric case where twists are always ‘global’ in some sense. Larger values of $\alpha_{*}^{\rm max}$ lead to more twisted configurations as expected, since the currents are stronger in this case. For example, comparing the final column we see a ‘weaving’ pattern only in in the larger twist case and that the toroidal blowout is a factor $\lesssim 2$ more effective.
@@
In Figure 5 we show the relative excess energy $E_{e}$ as a function of $\alpha_{*}^{\mathrm{max}}$. Interestingly, neither the functional dependence of the energy on $\alpha_{*}^{\rm max}$ nor the maximum achievable twist seem to be affected by the compactness. In particular, though a one-to-one comparison is made difficult by the different choices of current function, these features were also observed in stable axisymmetric cases (Fig. 3). Quantitatively however, because the twist is more localised the excess energy is lower. Note, however, that models with higher $M/R$ have higher absolute energy for a fixed dipole moment because the surface field is effectively amplified (Fig. 1).
@@ 5.1 Spindown luminosity and surface-field inferences
where $\Omega_{*}$ is the star-frame velocity. These facts together imply that the spindown luminosity, $L$, increases with compactness, as established by Rezzolla & Ahmedov (2004). In particular, since $L\propto B_{*}^{2}\Omega_{*}^{4}$ there is an effective amplification relative to the Newtonian luminosity by a factor $f_{R}^{2}/N_{R}^{4}$, where $f_{R}$ is the right-hand side of expression (13) and $N_{R}=e^{\nu(R)}=\Omega/\Omega_{*}$ (see equation 150 in Rezzolla & Ahmedov, 2004):
$\frac{L^{\rm GR}}{L^{\rm N}}=\kappa=f_{R}^{2}/N_{R}^{4}.$ (15)
This ratio, $\kappa$, scales sharply with compactness. For example, $\kappa=4.2$ for $M/R=0.17$ but $\kappa=10.7$ for $M/R=0.25$ (see also Sec. 5.3).
The equation obtained by matching $L^{\rm GR}$ to the rotational kinetic-energy loss can be rearranged to produce an expression for the (untwisted) polar field strength. Further correcting for the fact the star ought to be an oblique rotator surrounded by magnetospheric plasma rather than in vacuum (Pétri, 2016), we have (restoring units)
$B_{p}^{2}$ $=\frac{1}{\kappa\left(a+b\sin^{2}\chi\right)}\times\frac{3c^{3}P\dot{P}I_{0}}{2\pi^{2}R^{6}}$ (16)
$\approx\left(1-\frac{2GM}{c^{2}R}\right)^{7/2}\frac{3c^{3}P\dot{P}I_{0}}{2\pi^{2}R^{6}\left(a+b\sin^{2}\chi\right)},$
which one may recognize as the Spitkovsky (2006) formula multiplied by a redshift prefactor for (GR-corrected) moment of inertia $I_{0}$ and magnetic inclination angle $\chi$. The second line in equation (16) is simpler to use in practice and valid to within $3\%$ for $M/R\leq 0.2$. The factors $a$ and $b$ are fitting coefficients of order unity; in fact they are also generally larger than their Newtonian counterparts (see Table 2 in Pétri, 2016) indicating a further reduction. In the following, we take $a=1$ and $b=1.5$ and use the GR Tolman-VII moment of inertia, $I_{0}\approx 2MR^{2}(1-1.1M/R-0.6M^{2}/R^{2})^{-1}/7$ (Lattimer & Prakash, 2001), which is larger than the Newtonian value ($I_{0}^{N}=2MR^{2}/7$).
@@
$B^{\rm GR}_{p}\sim 1.6\times 10^{19}\sqrt{P\dot{P}}\text{ G},$ (17)
@@
Note that expression (17) should apply to the whole neutron-star population and not just magnetars.
Figure 6 displays some relevant dimensionless ratios, as functions of compactness, assuming the factors $a$ and $b$ in expression (16) are independent of $M/R$ and spin for simplicity (cf. Ruiz et al., 2014; Pétri, 2016). The top (dotted) line shows the results for the ratio of GR to Newtonian dipolar magnetospheric energy, assuming a fixed dipole moment $\mu_{B}$. The energy is obtained using equation (31). For example, for a star with $M/R=0.2$ the total energy is roughly doubled because the poloidal flux $\mathcal{P}$ is larger than the Newtonian value by a factor $\approx 1.4$ (see Fig. 1). In contrast, the bottom (dashed) line shows the same ratio, but for cases where the surface field is normalized under the assumption of a fixed $P\dot{P}$. The ratio $B_{p}^{GR}/B_{p}^{N}$ is deduced from equation (16) and is depicted by the solid line in the same Figure.
We see that while the energy itself scales with compactness (as per the dotted line), it does so at a rate that is shallower than the spindown luminosity (15). This implies that magnetar magnetospheres house less energy in the GR case if the field strength is deduced from timing parameters rather than assumed. For a canonical compactness of $M/R=0.17$, for example, the total magnetospheric energy is $\approx 38\%$ of the Newtonian value for any given $P\dot{P}$ (dashed line). The key conclusion is that since we find that the available free energy is almost constant as a function of compactness (see Figs. 3 and 5), a decrease by a factor $\sim 3$ in the (timing-inferred) energy available to fuel outbursts is expected for a typical neutron star ($M/R\sim 0.17$).
Figure 6: Ratios of dipolar magnetospheric energies normalised by the Newtonian values using the (GR or Newtonian) Tolman-VII equation of state as a function of compactness, with either fixed $\mu_{B}$ (dotted) or $P\dot{P}$ (dashed). The latter applies to cases where the field is inferred from timing data, with the ratio of surface-field inferences depicted by the solid line (see text).